OpenStudy (anonymous):

One more problem to check please ^_^ , again will take me a second to type all

OpenStudy (anonymous):

\[9^{x^2-2x}=27^{x^2+1}\] My final answer was x=-3 x=-1, both were invalid when I plugged them back into the equation. Is that possible?

OpenStudy (johnweldon1993):

Well lets do it this way...how did you arrive at your results?

OpenStudy (johnweldon1993):

And by that I mean your result of "Plugging them back in giving an invalid result"

OpenStudy (johnweldon1993):

Because when I plug them in I certainly get the correct result

OpenStudy (anonymous):

\[(3^2)^{x^2-2x}=(3^3)^{x^2+1}\] \[\log_{3}3^{2x^2-4x}=\log_{3}3^{3x^2+3} \] \[2x^2-4x=3x^2+3\] \[x^2+4x+3=0\]

OpenStudy (anonymous):

X=-3,-1

OpenStudy (johnweldon1993):

Right...-3 and -1 are indeed the correct roots...but now show the steps of plugging them back in to check

OpenStudy (anonymous):

and then i plugged back in the values of x back into the original equation, but they didn't seem to add up when i did so

OpenStudy (phi):

-1 definitely works.

OpenStudy (anonymous):

i see what i was doing wrong

OpenStudy (anonymous):

i messed up my negatives on my calculator; wasn't isolating them properly.

OpenStudy (johnweldon1993):

Ahh...that'll do it :)

OpenStudy (anonymous):

thank you